OpenStudy (anonymous):

More elimination! (NO clue how to do this one) 4x-y=-24 3x-y=-19

OpenStudy (mertsj):

1. Multiply the second equation by -1. 2. Add that result to first equation.

OpenStudy (anonymous):

Ok I wasn't sure if it was -1.

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

-3x+y=19?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

x=-5

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

Now find y by substitution.

OpenStudy (anonymous):

y=-4

OpenStudy (mertsj):

Better check your signs. There's an error somewhere.

OpenStudy (mertsj):

\[4(-5)-y=-24\] \[-20-y=-24\] \[-y=-24+20\] \[-y=-4\] \[y=4\]

OpenStudy (anonymous):

4(-5)-y=-24 -20-y=-24 +20 +20 y=-4 ok i disregarded the extra - in front of the y

OpenStudy (mertsj):

yes...can't do that.

OpenStudy (anonymous):

hahaha yeahhhhh dangit. Thank you!